OpenStudy (anonymous):

I need your help on what to do and how to start , Hockey team receives 2 point for a win 1 point for a tie. The wildcats once won a championship with 60 points. They won 9 more games than they tied. How many wins and how many ties did the wildcats have?

OpenStudy (anonymous):

w = number of games won t = number of games tied w = t+9 points: 2*w+1*t = 60

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

So you will now start with 2*(t+9) + t = 60 2t + 18 + t = 60 3t = 42 t = 14

OpenStudy (anonymous):

and then you will finish with w = t + 9 w = 14 + 9 w = 23. So you will have 23 wins and 14 ties.

OpenStudy (anonymous):

so I would basically be using the order of operations or a different formula ?